OpenStudy (owlcoffee):

Is my proof correct? (0=1)?

OpenStudy (owlcoffee):

\[0=0+0+0+0+...\] \[0=(1-1)+(1-1)+(1-1)+...\] \[0=1-1+1-1+1-1+...\] \[0=1+(-1+1)+(-1+1)+(-1+1)+...\] Therefore: \[0=1\]

Parth (parthkohli):

no

OpenStudy (owlcoffee):

Well, why?

Parth (parthkohli):

any proof that gets the wrong result isn't correct. :P

Parth (parthkohli):

From what I see, it's not the best way to handle a divergent series.

OpenStudy (mayankdevnani):

hahaha.... its absolutely wrong

random231 (random231):

is this a joke?:D

random231 (random231):

wasting threads!!

OpenStudy (owlcoffee):

Well, I see no justifications, so I don't think so.

OpenStudy (mayankdevnani):

see, if we have odd like 5 ones :- \[\huge \bf 0=1-1+1-1+1\] \[\large \bf 0=1+(-1+1)+(-1+1)=\frac{0}{2}+0+0\]

OpenStudy (mayankdevnani):

is it correct ? @Owlcoffee

Parth (parthkohli):

You're bringing a series that diverges, i.e., has no value.

random231 (random231):

x = y. Then x2 = xy. Subtract the same thing from both sides: x2 - y2 = xy - y2. Dividing by (x-y), obtain x + y = y. Since x = y, we see that 2 y = y. Thus 2 = 1, since we started with y nonzero. Subtracting 1 from both sides, 1 = 0.

random231 (random231):

hue hue hue

OpenStudy (mayankdevnani):

hope you understand

OpenStudy (owlcoffee):

Wow, I don't want to see you guys to write a proof, ever. I want solid arguements, not arrows and "Oh, duh, he's dumb".

random231 (random231):

he kinda sent the 1st 1 to the end of the series,,, :D

random231 (random231):

u cannot write a series like dat!

OpenStudy (hihi67):

Dude between step 2 and 3, you can't just take the brackets away.

Parth (parthkohli):

You start off with the supposition that (1 - 1) + (1 - 1) ... = 1 - 1 + 1 - 1 ... This is an infinite series, and you can't change the order of computation.

OpenStudy (hihi67):

^

OpenStudy (mayankdevnani):

you want solid arguments :- here :- first, \[\large \bf 0=0+0+0+0+0\] \[\large \bf 0=(1-1)+(1-1)+(1-1)+(1-1)+(1-1)\] open brackets, \[\large \bf 0=1-1+1-1+1-1+1-1+1-1\] then, arrange it \[\bf 0=1+(-1+1)+(-1+1)+(-1+1)+(-1+1)-1\] \[\large \bf 0=1-1\] \[\large \bf 0=0\] Hence, Proved.

OpenStudy (akashdeepdeb):

Let us assume 0 = 2 0 * 0 = 2 * 0 = 0 Thus, both of them are same? WE CANNOT SAY THAT! :D

OpenStudy (mayankdevnani):

hope you now understand. @Owlcoffee

OpenStudy (akashdeepdeb):

Although if you want to a cool math trick, which is actually legitimate and is used with lorentzian transformations in spacetime events. 1 + 2 + 3 + 4 + 5..... = -1/12 Google it. :)

OpenStudy (mayankdevnani):

i have proved it. @AkashdeepDeb

OpenStudy (mayankdevnani):

in the style of owlcoffee

OpenStudy (aravindg):

http://mathworld.wolfram.com/DivergentSeries.html

OpenStudy (owlcoffee):

Sweg

OpenStudy (mayankdevnani):

Satisfied ? @Owlcoffee

OpenStudy (owlcoffee):

Thank you. I knew how it was all along, ir was just a way for me to evaluate how people view mathematics. It's a group proyect I have.